# RHanDS: Refining Malformed Hands for Generated Images with Decoupled Structure and Style Guidance

### v1

@@ Abstract.
Although diffusion models can generate high-quality human images, their applications are limited by the instability in generating hands with correct structures. Some previous works mitigate the problem by considering hand structure yet struggle to maintain style consistency between refined malformed hands and other image regions. In this paper, we aim to solve the problem of inconsistency regarding hand structure and style. We propose a conditional diffusion-based framework RHanDS to refine the hand region with the help of decoupled structure and style guidance. Specifically, the structure guidance is the hand mesh reconstructed from the malformed hand, serving to correct the hand structure. The style guidance is a hand image, e.g., the malformed hand itself, and is employed to furnish the style reference for hand refining. In order to suppress the structure leakage when referencing hand style and effectively utilize hand data to improve the capability of the model, we build a multi-style hand dataset and introduce a two-stage training strategy. In the first stage, we use paired hand images for training to generate hands with the same style as the reference. In the second stage, various hand images generated based on the human mesh are used for training to enable the model to gain control over the hand structure. We evaluate our method and counterparts on the test dataset of the proposed multi-style hand dataset. The experimental results show that RHanDS can effectively refine hands structure- and style- correctly compared with previous methods. The codes and datasets will be available soon.
malformed hand refining, diffusion models, conditional generation
copyright: none
@@
Figure 1. Examples of hands refined by our RHanDS (right in each pair) from malformed hand (left in each pair). Typical malformations of generated hands include irregular hand shapes and incorrect numbers of fingers.
@@
Figure 2. The RHanDS framework we propose contains four modules: a VAE for projecting images into a latent space and reconstructing images from the latent, a conditional U-net for predicting the denoised variant during the denoising process, a style encoder to extract hand style from the malformed hand and map it into the U-net via cross-attention, and a structure encoder to utilize the hand mesh reconstructed from the malformed hand to guide the hand structure. In addition, to achieve a fully automatic process, a hand detection model and a 3D hand reconstruction model are required.
1. Introduction
@@
Text-to-image diffusion models (Nichol et al., 2022; Rombach et al., 2022; Ramesh et al., 2022; Saharia et al., 2022; Podell et al., 2023) have exhibited the remarkable ability to synthesize visually stunning images based on textual prompts, representing a significant advancement in image generation. Despite their impressive capabilities, current models still face challenges in effectively handling intricate structures like hands (Podell et al., 2023; Rombach et al., 2022). As illustrated in Figure 1, models often produce malformed hands with irregular shapes or incorrect numbers of fingers, deviating from the realistic 3D shape and physical limitations of human hands.
To solve the problem, Lu et al. (Lu et al., 2023) introduces a post-processing solution that employs an inpainting diffusion model to refine the malformed hand in the generated image. During the inpainting process, a 3D hand mesh is reconstructed based on the malformed hand image, rendered as a depth image to provide structural guidance for correcting the malformed hand through ControlNet (Zhang et al., 2023). This inpainting method, which incorporates hand structure control, indeed effectively enhances the structural accuracy of generating hand images. However, the generated hand image may lack consistency in color, texture, and other stylistic features compared to the original malformed hands, resulting in a lack of harmony with the overall image. This issue will become more serious when dealing with hand styles that are less common in datasets, such as in sculpture or cartoon images. It may be due to the indirect access to hand style information, requiring implicit inference of hand texture and color from other skin areas in the image during the inpainting process. In order to alleviate this phenomenon, it is intuitive to consider utilizing the image region of the original malformed hand to enhance style consistency. However, it remains an open challenge to selectively incorporate hand style features without introducing structural information (Ye et al., 2023b), in order to avoid interference with the correct hand structure.
@@
In this paper, we propose a novel framework named RHanDS to Refine malformed Hands with Decoupled Structure and Style guidance, which is adapted to malformed hands of arbitrary styles. As shown in Figure 2, RHandDS repaints the hand region based on diffusion models to refine the malformed hand structure while preserving the hand style. Specifically, RHandDS utilizes the 3D mesh reconstructed from the malformed hand for structural guidance via the rendered depth image and structure encoder, and uses the style encoder to extract the style embedding from the malformed hand for style guidance.
@@
In order to suppress the leakage of hand structure from style embedding, we also introduce a two-stage training strategy and propose corresponding hand datasets. The purpose of the first-stage training is to equip the model with the capability of repainting hands with a given style. In order to extract hand styles while ignoring structural information, we collect a large number of realistic image pairs depicting the left and right hands of the same person (i.e., images that share the same style but exhibit different structures). We then use one hand as a style reference to guide the generation of the other hand. In this stage, we only train the style encoder and U-net, with the structure encoder remaining unutilized. In the second stage, our objective shifts towards making the model generate hand images with additional structural control. Considering the limited style diversity in existing hand image datasets (with 3D meshes, data like these are called as hand-mesh pairs for convenience in this paper), we propose a novel approach using the SMPL model and ControlNet to create a more diverse dataset. We initiate this process by generating hand meshes using the SMPL model and rendering the corresponding depth images. ControlNet is then utilized to generate hand images, controlled with depth images and prompts indicating various style categories. The generated hand images are evaluated by calculating the Mean Per Joint Position Error (MPJPE) (Ionescu et al., 2013) metric. Images with small MPJPE differences are deemed to have higher structural accuracy and are retained for training the structure encoder while maintaining the network’s other parameters unchanged.
@@
Besides, to verify the hand refining ability, we use a standard diffusion model to regenerate the hand regions in the HumanArt dataset (Ju et al., 2023), and propose a multi-style malformed hand dataset with mesh-image pairs for evaluation. Qualitative and quantitative experiments show the effectiveness and superiority of RHanDS.
@@
  • For refining malformed hand images, we propose a novel framework named RHandDS with decoupled structure and style guidance. Benefiting from direct and decoupled style guidance for malformed hands, our model can maintain stylistic consistency before and after hand refining and generate hand images in diverse styles.
@@
  • We propose a two-stage training strategy and create corresponding datasets, which effectively enable the model to learn control over hand style and hand structure separately.
@@
  • Experiments show that RHanDS can effectively handle malformed hands in various styles and restore them to better structures.
2. Related Work
2.1. Diffusion Models for Image Synthesis
@@
Recent advancements have been greatly promoted by text-to-image diffusion models. Specifically, the technique of Latent Diffusion Models (LDM) (Rombach et al., 2022) conducts diffusion in a latent image space (Van Den Oord et al., 2017), which largely reduces computational demands. Stable Diffusion (AI, 2022b) is one of the highlighted works of LDM, which utilizes a pre-trained language encoder like CLIP (Radford et al., 2021) to encode the text prompt into latent space to guide the diffusion process.
@@
Diverging from the text-to-image paradigm, image inpainting is tasked with synthesizing visually plausible content within the masked regions of the existing image, ensuring that the synthesized content maintains semantic coherence with the context of the unmasked regions. By fine-tuning the pre-trained text-to-image Stable Diffusion model, the specialized variant known as Stable Diffusion Inpainting (AI, 2022a) can synthesize content within the masked region that follows textual prompts. However, fine-grained control over the structure and style of the inpainting contents cannot be achieved solely through text guidance.
@@
In order to provide fine-grained conditions during the generation process, different control methods have been explored in previous studies, and it has been demonstrated that an additional module can be effectively plugged into the existing diffusion models to guide the image generation. ControlNet (Zhang et al., 2023) and T2I-Adapter (Mou et al., 2024) use spatial conditions such as canny, depth images, and body poses as inputs to control the structure of the generated image. In order to reduce the fine-tuning cost, Uni-ControlNet (Zhao et al., 2024) allows for the simultaneous utilization of different conditions within one single model. Previous subject-driven approaches such as DreamBooth (Ruiz et al., 2023), Textual Inversion (Gal et al., 2022), LoRA (Hu et al., 2021), and Concept Sliders (Gandikota et al., 2023) achieved stylization and customization by fine-tuning parameters. In contrast, IP-Adapter (Ye et al., 2023b) leverages pre-trained image encoder (Radford et al., 2021) to achieve the image prompt capability for diffusion models, generating images that resemble the reference images in content and style.
2.2. Plausible Hand Generation
@@
Concept Sliders (Gandikota et al., 2023) use parameter-efficient training method (Hu et al., 2021) to learn better physical structure of hand by fine-tuning diffusion models with carefully selected data. Yeet al. (Ye et al., 2023a) synthesize hand-object interaction on existing images using the rough region of the palm and forearm as guidance. However, the inherent complex structure of 3D human hands makes it difficult for these methods to generate correct hands stably. Different from the methods that directly generate hands, some works (Ye et al., 2023a; Lu et al., 2023; Weng et al., 2023) introduce the structure information of the target hand into a model to prevent the model from generating malformed hands. HanDiffuser (Ye et al., 2023a) encodes the 3D hand parameters into text embedding for achieving text-to-image generation with realistic hand appearances. Diffusion-HPC (Weng et al., 2023) proposes a post-processing method that refines the generated malformed person through a conditional diffusion model with the help of a human depth image rendered based on the reconstructed human mesh. HandRefiner (Lu et al., 2023) proposes a similar post-processing method more suitable for dealing with malformed hands.
@@
Our RHanDS is a post-processing method that leverages 3D hand mesh as the pixel-level condition to control hand structure. Compared to existing methods, RHanDS performs specifically on the hand region rather than the entire image, which facilitates more precise refinement. Moreover, RHanDS enhances the perception of hand style, ensuring that the style of the refined hand seamlessly aligns with that of the original image for a coherent and authentic representation.
3. methodology
@@
Given a generated image that contains malformed human hands, we aim to correct the structure of the malformed hand. To this end, in this section, we introduce the detail of our proposed diffusion-based framework RHanDS. As shown in Figure 2, RHanDS mainly contains four modules: style encoder $\mathcal{E}_{style}$, structure encoder $\mathcal{E}_{struc}$, U-net $\epsilon_{\theta}$ and VAE $\mathcal{E}$ and $\mathcal{D}$. Specifically, the inputs of the whole framework are a cropped malformed hand image $x$, a structure reference $ref_{struc}$, a style reference $r_{style}$, and a mask $mask$. The malformed hand image $x$ is cropped out from the entire image for effective refining. The structure reference $ref_{struc}$ is a hand depth image rendered from a hand model that can be manually created using 3D tools (ZhUyU1997, [n. d.]) or automatically obtained from the 3D mesh reconstructed from malformed hand and is encoded by the structure encoder $\mathcal{E}_{struc}$ to guide the hand structure. The style reference $r_{style}$ is a hand image that can either be the malformed hand itself or any other hand image and is encoded by the style encoder $\mathcal{E}_{style}$ to guide the hand style. The binary mask $mask$ indicates the malformed hand region that needs to be repainted. The overall training process is decomposed into two stages to decouple the structure and style guidance. In the first stage (Section 3.2), U-net and style encoder are trained using multi-style paired hand images. In the second stage (Section 3.3), the structure encoder is trained using the multi-style hand-mesh dataset.
Our RHanDS ensures the correctness of the structure with a 3D hand model, maintains the correctness of the style with reference hands, and avoids the mutual influence that exists between structure and style guidance. As a result, RHanDS can produce plausible hand images $x^{\prime}$ with the correct structure while preserving the hand style.
3.1. Preliminaries
@@
RHanDS is built on top of stable diffusion, which performs the diffusion process efficiently in the latent space rather than pixel space. Specifically, given an image $x_{0}$ in RGB space, the encoder $\mathcal{E}$ first encodes the image into a latent representation $z_{0}=\mathcal{E}(x_{0})$. During the forward process, normally-distributed noise $\epsilon_{t}$ is added into the latent $z_{0}$ to obtain the noisy latent $z_{t}$. During the reverse process, the stable diffusion model implements denoising U-net $\epsilon_{\theta}$ as the backbone to predict the noise $\epsilon_{t}$ with noisy latent $z_{t}$ and current timestep $t$. The simple training objective can be written as
(1) $\mathcal{L}=\mathbb{E}_{z_{0},\epsilon\sim\mathcal{N}(0,1),t}[||\epsilon-\epsilon_{\theta}(z_{t},t)||^{2}_{2}].$
@@
To correct the structure of malformed hand in image while keeping other region unchanged, the underlying neural backbone of RHanDS is implemented as the U-net of stable diffusion inpainting model, a particular variant of stable diffusion, which enables the functionality of inpainting specified regions of an image under the guidance of a mask. Compared to the standard version of stable diffusion, this U-net has five additional input channels for supporting inpainting. Specifically, four of these channels are allocated for the encoded masked-image $\hat{z}_{0}$, while the remaining single channel is allocated for the resized mask $mask$. During reverse process, the U-net predicts the noise through $\epsilon_{t}=\epsilon_{\theta}(z_{t},\hat{z}_{0},mask,t)$ while $\hat{z}_{0}$ and $mask$ keep unchanged. After the reverse process, the decoder $\mathcal{D}$ reconstructs the image from the latent.
Figure 3. The first-stage training of RHanDS. In this stage, U-net and style encoder are trained using multi-style paired hand images for style guidance.
3.2. The First Stage: Style Guidance
As shown in Figure 6, the existing solutions are limited in the perception of hand style and exhibit poor generalization ability in images with anime, oil painting, watercolor, and other styles. To solve the problem, we have to encode the hand style into U-net, enabling the model to refine hands $x$ based on style reference $r_{style}$. Similar to the approach used in IP-Adapter (Ye et al., 2023b), we use a CLIP (Radford et al., 2021) image encoder to extract the image embedding from $r_{style}$. To enforce the CLIP model extract styles from hand images without introducing other extraneous information such as hand structure, we choose the global image embedding and use a linear projection network to transform the image embedding into a sequence of features with length $tokens$ to obtain the style embedding $c_{style}$. Overall, the style encoder $\mathcal{E}_{style}$ consists of a CLIP image encoder and a linear projection network, and style embedding $c_{style}$ that can be formulated as follows:
(2) $c_{style}=\mathcal{E}_{style}(r_{style})$
As the style embedding already contains sufficient information for hand refining, we drop the text features and directly feed the style embedding into U-net through cross-attention layers. Given the style embedding, the output of cross-attention $\mathbf{Z}$ is computed as follows:
(3) $\mathbf{Z}=Attention(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{Softmax}(\frac{\mathbf{Q}\mathbf{K^{T}}}{\sqrt{d}})\mathbf{V},$
where $\mathbf{Q}=\mathbf{Z}\mathbf{W}_{q},\mathbf{K}=c_{style}\mathbf{W}_{k},\mathbf{V}=c_{style}\mathbf{W}_{v}$ are query, key, value matrices of the attention layer respectively, and $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, $\mathbf{W}_{v}$ are the weight matrices of the linear projection layers in attention module.
As shown in Figure 3, in this stage, we train the U-net and the style encoder on the dataset with multi-style paired hands, using the same training objective as stable diffusion:
(4) $\mathcal{L}=\mathbb{E}_{z_{t},\epsilon,t}[||\epsilon-\epsilon_{\theta}(z_{t},c_{style},mask,t)||^{2}_{2}].$
where $c_{style}$ used to supply style guidance is encoded from one hand image, and $z_{0}$ used for training U-net to predict noise is encoded from the other hand image $x$. It is worth noting that $r_{style}$ needs to be different from $x$ in terms of gesture while maintaining consistency in hand style, and the $r_{style}$ and $x$ must exhibit diversity in styles. Otherwise, the style encoder $\mathcal{E}_{style}$ would encode the gesture of the $r_{style}$ for noise prediction or fail to encode style guidance for U-net. To build the multi-style paired hand dataset that meets the requirements, we crop the two hands of the same character that exhibit the same style but different gesture from the characters images of varying styles. The details of this dataset can be found in Subsection 4.1.
Figure 4. The second-stage training of RHanDS. In this stage, structure encoder is trained using multi-style hand-mesh data for structure guidance.
3.3. The Second Stage: Structure Guidance
With the assistance of the style encoder, our framework can utilize the style reference $r_{style}$ to guide the refining process of malformed hands. However, as shown in Figure 6, it is hard to correct the structure of the malformed hand stably due to the inherent complex structure of 3D human hands. To solve the problem, we adopt MANO (Romero et al., 2017a), a 3D hand model with articulated and non-rigid deformations, to represent the correct gesture and structure of hands. Specifically, we used MANO-based methods to reconstruct the 3D hand mesh from the malformed hand image. For malformed hands that are difficult to reconstruct automatically, this process can also be completed manually using existing tools (ZhUyU1997, [n. d.]). The correct hand structure has been determined by the 3D hand mesh, and following (Lu et al., 2023), we render the hand depth image from the hand mesh as structure reference $r_{struc}$ to guide the correction of hand structure.
To encode the structure reference $r_{struc}$ into U-net for structure guidance, we use the structure encoder $\mathcal{E}_{struc}$ to encode the structure reference $r_{struc}$, and $\mathcal{E}_{struc}$ adopts the same structure as ControlNet (Zhang et al., 2023). Different from the original setting in (Zhang et al., 2023), we feed a learnable embedding $c_{l}$ into the cross-attention layers of $\mathcal{E}_{struc}$ to maintain the performance of pre-trained model. Specially, the input of $\mathcal{E}_{struc}$ is consist of structure reference $r_{struc}$, noisy latent $z_{t}$ and learnable embedding $c_{l}$. The output of $\mathcal{E}_{struc}$ is the structure embedding $c_{struc}$ that can be formulated as follows:
(5) $c_{struc}=\mathcal{E}_{struc}(r_{struc},z_{t},c_{l})$
For all the outputs $out^{i}$ of the middle block and decoder blocks of U-net, the structure embedding $c_{struc}^{i}$ is fed into U-net through a $1\times 1$ convolution layer $\mathbb{Z}^{i}$ with structure weight $w$:
(6) $out^{i}=out^{i}+w\cdot\mathbb{Z}^{i}(c_{struc}^{i})$
As shown in Figure 4, in this stage, we freeze the U-net and the style encoder to maintain the capability of style guidance, and train the structure encoder on the multi-style hand-mesh dataset, using the training objective as follows:
(7) $\mathcal{L}=\mathbb{E}_{z_{t},\epsilon,t}[||\epsilon-\epsilon_{\theta}(z_{t},c_{style},c_{struc},mask,t)||^{2}_{2}].$
where $c_{style}$ and $z_{0}$ are encoded from the same hand image, and $c_{struc}$ is encoded from the hand depth image obtained by rendering the 3D hand mesh. However, the model tends to overfit the hand styles present in the dataset due to style limitations inherent within the existing hand-mesh dataset (AI., [n. d.]; Zimmermann and Brox, 2017; Chen et al., 2022). A potential solution is to employ a depth estimation model (Bhat et al., 2023) to predict hand depth from multi-style hand datasets, but the noise inherent in the depth prediction may undermine the capability to control hand structure. Based on the depth-controlled text-to-image generation pipeline, we use stable diffusion and ControlNet to create a multi-style hand-mesh dataset by using the depth image rendered from random hand meshes. The details of this dataset can be found in Section 4.2.
Figure 5. The multi-style datasets. Multi-Style Paired Hand Dataset includes hand images of both hands of characters. Multi-Style Hand-Mesh Dataset includes hand meshes and hand images generated based on meshes. Multi-Style Malformed Hand-Mesh Dataset includes real hands, malformed hands generated from real hands, and hand meshes reconstructed from real hands.

### v2

@@ Abstract
Although diffusion models can generate high-quality human images, their applications are limited by the instability in generating hands with correct structures. In this paper, we introduce RHanDS, a conditional diffusion-based framework designed to refine malformed hands by utilizing decoupled structure and style guidance. The hand mesh reconstructed from the malformed hand offers structure guidance for correcting the structure of the hand, while the malformed hand itself provides style guidance for preserving the style of the hand. To alleviate the mutual interference between style and structure guidance, we introduce a two-stage training strategy and build a series of multi-style hand datasets. In the first stage, we use paired hand images for training to ensure stylistic consistency in hand refining. In the second stage, various hand images generated based on human meshes are used for training, enabling the model to gain control over the hand structure. Experimental results demonstrate that RHanDS can effectively refine hand structure while preserving consistency in hand style.
Codehttps://github.com/alimama-creative/RHanDS
@@
Figure 1: Examples of hands refined by our RHanDS (right in each pair) from the malformed hands (left in each pair).
@@
Figure 2: The RHanDS framework we propose contains four modules: a VAE for projecting images into a latent space and reconstructing images from the latent, a conditional U-net for predicting the denoised variant during the denoising process, a style encoder to extract hand style from the malformed hand and map it into the U-net via cross-attention, and a structure encoder to utilize the hand mesh reconstructed from the malformed hand to guide the hand structure. In addition, to achieve a fully automatic process, a hand detection model and a 3D hand reconstruction model are required.
1 Introduction
@@
Text-to-image diffusion models (Nichol et al. 2022; Rombach et al. 2022; Ramesh et al. 2022; Saharia et al. 2022; Podell et al. 2023) have exhibited the remarkable ability to synthesize visually stunning images based on textual prompts, representing a significant advancement in image generation. Despite their impressive capabilities, current models still face challenges in effectively handling intricate structures like hands (Podell et al. 2023; Rombach et al. 2022). As illustrated in Figure 1, models often produce malformed hands with irregular shapes or incorrect numbers of fingers, deviating from the realistic 3D shape and physical limitations of human hands.
To address the problem, Lu et al. introduces a post-processing framework that employs an inpainting diffusion model to refine the malformed hands. During the refining process, a 3D hand mesh is reconstructed based on the malformed hand image and rendered into a depth image to provide structure guidance for correcting the malformed hand through ControlNet (Zhang, Rao, and Agrawala 2023). However, it is difficult to maintain the hand style as the refining process requires adding noise to the hand image. This issue becomes particularly serious when addressing specific styles like sculptures or cartoons. To enhance style consistency, it seems intuitive to utilize the image region of the original malformed hand. However, extracting style information from the malformed hand inevitably introduces incorrect structural information (Ye et al. 2023a).
@@
In this paper, we propose a novel framework named RHanDS to refine malformed Hands with Decoupled Structure and Style guidance, which is adaptable to malformed hands of arbitrary styles. As shown in Figure 2, RHandDS utilizes diffusion models to repaint the hand region, effectively refining malformed hand structures while preserving the original hand style. In detail, the 3D mesh reconstructed from the deformed hand is rendered into a depth image, which is then input into the structure encoder to provide structure guidance. Meanwhile, the style encoder extracts the style embedding from the malformed hand to offer style guidance. To alleviate the mutual interference between style and structure guidance, we introduce a two-stage training strategy and propose the corresponding multi-style hand datasets.
The first stage focuses on enabling the model to repaint hands in a specified style. To isolate hand styles from structural information, we collect a large number of realistic image pairs depicting the left and right hands of the same individual (i.e., images that share the same style but exhibit different structures). We then use one hand as a style reference to guide the generation of the other hand. In this stage, we only train the style encoder and U-net.
@@
In the second stage, our objective shifts towards making the model generate hand images with additional structural control. Considering the limited style diversity in existing hand image datasets (with 3D meshes, data like these are called hand-mesh pairs for convenience in this paper), we propose a novel approach using the SMPL model and ControlNet to create a more diverse dataset. We initiate this process by generating hand meshes using the SMPL-H model and rendering the corresponding depth images. ControlNet is then utilized to generate hand images, controlled with depth images and prompts indicating various style categories. In this stage, we train the structure encoder while keeping the other parameters frozen.
@@
To verify the ability of hand refining, we use a standard diffusion model to regenerate the hand regions in the HumanArt dataset (Ju et al. 2023), and propose a multi-style malformed hand dataset with mesh-image pairs for evaluation. Both qualitative and quantitative experiments demonstrate the effectiveness and superiority of RHanDS.
@@
  • To refine malformed hand images, we propose a novel framework named RHandDS, which adopts decoupled structure and style guidance. By leveraging the direct and decoupled style guidance, our model can consistently maintains hand style while refining hand images across various styles.
@@
  • We propose a two-stage training strategy and create corresponding datasets, enabling the model to learn control over hand style and structure separately.
@@
  • Experiments demonstrate that RHanDS can effectively handle malformed hands across various styles and restore them to better structures.
2 Related Work
2.1 Diffusion Models for Image Synthesis
@@
Recent advancements have been greatly promoted by text-to-image diffusion models. Specifically, the technique of Latent Diffusion Models (LDM) (Rombach et al. 2022) conducts diffusion in a latent image space (Van Den Oord, Vinyals et al. 2017), which largely reduces computational demands. Stable Diffusion (StabilityAI 2022b) is one of the highlighted works of LDM, which utilizes a pre-trained language encoder like CLIP (Radford et al. 2021) to encode the text prompt into latent space to guide the diffusion process.
@@
Diverging from the text-to-image paradigm, image inpainting is tasked with synthesizing visually plausible content within the masked regions of the existing image, ensuring that the synthesized content maintains semantic coherence with the context of the unmasked regions. By fine-tuning the pre-trained text-to-image Stable Diffusion model, the specialized variant known as Stable Diffusion Inpainting model (StabilityAI 2022a) can synthesize content within the masked region that follows textual prompts. However, fine-grained control over the structure and style of the inpainting contents cannot be achieved solely through text guidance.
@@
In order to provide fine-grained conditions during the generation process, different control methods have been explored in previous studies, and it has been demonstrated that an additional module can be effectively plugged into the existing diffusion models to guide the image generation. ControlNet (Zhang, Rao, and Agrawala 2023) and T2I-Adapter (Mou et al. 2024) use spatial conditions such as canny, depth images, and body poses to control the structure of the generated image. To reduce the fine-tuning cost, Uni-ControlNet (Zhao et al. 2024) allows for the simultaneous utilization of different conditions within one single model. Previous subject-driven approaches such as DreamBooth (Ruiz et al. 2023), Textual Inversion (Gal et al. 2022), LoRA (Hu et al. 2021), and Concept Sliders (Gandikota et al. 2025) achieved stylization and customization by fine-tuning parameters. In contrast, IP-Adapter (Ye et al. 2023a) leverages pre-trained image encoder (Radford et al. 2021) to achieve the image prompt capability for diffusion models, generating images that resemble the reference images in content and style.
2.2 Plausible Hand Generation
@@
Concept Sliders (Gandikota et al. 2025) use parameter-efficient training method (Hu et al. 2021) to learn better physical structure of hand by fine-tuning diffusion models with carefully selected data. Ye et al. synthesize hand-object interaction on existing images using the rough region of the palm and forearm as guidance. However, the inherent complex structure of 3D human hands makes it difficult for these methods to generate correct hands stably. Different from the methods that directly generate hands, some works (Ye et al. 2023b; Lu et al. 2024; Weng, Bravo-Sánchez, and Yeung 2023) introduce the structure information of the target hand into a model to prevent the model from generating malformed hands. HanDiffuser (Ye et al. 2023b) encodes the 3D hand parameters into text embedding for achieving text-to-image generation with realistic hand appearances. Diffusion-HPC (Weng, Bravo-Sánchez, and Yeung 2023) proposes a post-processing method that refines the generated malformed person through a conditional diffusion model with the help of a human depth image rendered based on the reconstructed human mesh. HandRefiner (Lu et al. 2024) proposes a similar post-processing method more suitable for dealing with malformed hands.
@@
Our RHanDS is a post-processing method that leverages 3D hand mesh as the pixel-level condition to control hand structure. Compared to existing methods, RHanDS performs specifically on the hand region rather than the entire image, facilitating more precise refinement. Moreover, RHanDS enhances the perception of hand style, ensuring that the style of the refined hand seamlessly aligns with that of the original image for a coherent and authentic representation.
3 Methodology
@@
In this section, we introduce the detail of our proposed framework RHanDS. Given a generated image containing malformed human hands, we aim to correct the structure of the malformed hand while preserving the style of the generated image. As shown in Figure 2, RHanDS mainly contains four modules: style encoder $\mathcal{E}_{style}$, structure encoder $\mathcal{E}_{struc}$, U-net $\epsilon_{\theta}$, and VAE encoder $\mathcal{E}$ and decoder $\mathcal{D}$. The inputs of the framework include a cropped malformed hand image $x$, a structure reference $r_{struc}$, a style reference $r_{style}$, and a mask $mask$. The malformed hand image $x$ is cropped out from the entire image for effective refining. The structure reference $r_{struc}$ is a depth image rendered from a 3D hand model. The 3D hand model can be manually created using 3D tools (ZhUyU1997 2023) or automatically reconstructed from malformed hand, and the structure reference is encoded by the structure encoder $\mathcal{E}_{struc}$ to guide the hand structure. The style reference $r_{style}$ is the malformed hand itself and is encoded by the style encoder $\mathcal{E}_{style}$ to guide the hand style. The binary mask $mask$ indicates the malformed hand region that needs to be repainted. The training process is decomposed into two stages to decouple the structure and style guidance. In the first stage, the U-net and style encoder are trained using Multi-Style Paired Hand Dataset. In the second stage, the structure encoder is trained using the Multi-Style Hand-Mesh Dataset.
3.1 Preliminaries
@@
RHanDS is built on stable diffusion, which efficiently performs the diffusion process in latent space rather than pixel space. Specifically, given an image $x_{0}$ in RGB space, the encoder $\mathcal{E}$ first encodes the image into a latent representation $z_{0}=\mathcal{E}(x_{0})$. During the forward process, normally-distributed noise $\epsilon_{t}$ is added into the latent $z_{0}$ to obtain the noisy latent $z_{t}$. During the reverse process, the stable diffusion model implements denoising U-net $\epsilon_{\theta}$ as the backbone to predict the noise $\epsilon_{t}$ with noisy latent $z_{t}$ and current timestep $t$. The simple loss function can be written as
$\mathcal{L}=\mathbb{E}_{z_{0},\epsilon\sim\mathcal{N}(0,1),t}[||\epsilon-\epsilon_{\theta}(z_{t},t)||^{2}_{2}].$ (1)
@@
In this work, RHanDS employs a U-net with five additional input channels. Specifically, four channels are allocated for the encoded masked-image $\hat{z}_{0}$, and the remaining single channel is allocated for the resized mask $mask$. During reverse process, the U-net predicts the noise through $\epsilon_{t}=\epsilon_{\theta}(z_{t},\hat{z}_{0},mask,t)$, where $\hat{z}_{0}$ and $mask$ keep unchanged. After the reverse process, the decoder $\mathcal{D}$ reconstructs the hand image from the latent.